OpenStudy (anonymous):

write an equation in the form y=mx+b for the line that has a slope of 1/4 and contains the point (12,3)

OpenStudy (nayef):

m=the slope To find b: substitute the slope and the point and then solve; like: 3=m(12)+b

OpenStudy (anonymous):

1/4 is the slope

OpenStudy (anonymous):

yes so that gives us y=1/4 +3

OpenStudy (anonymous):

because "b" is the y-intercept

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

said i was wrong

pooja195 (pooja195):

still need help?

OpenStudy (anonymous):

lol yes i do geting as much notes as i can

pooja195 (pooja195):

Just like the last one we need to start off with the formula \[\huge~y-y1=m(x-x1)\] 1/4=m 12=x1 3=y1

OpenStudy (anonymous):

the last one we did was wrong too for some reason

pooja195 (pooja195):

hmmm :/ \[\huge~y-3=1/4(x-12)\]

pooja195 (pooja195):

distribute 1/4 * x =? 1/4 * -12 = ?

OpenStudy (anonymous):

-3

pooja195 (pooja195):

\[\huge~y−3+3= 1 4 x−3+3 \]

pooja195 (pooja195):

we add 3 to both sides and then everything cancels and now we are only left with y=1/4x

pooja195 (pooja195):

^that should be correct